The decibel level of sound is given by D= 10log ( I/10^−12), where I is the
sound intensity measured in watts per square meter. Find the
decibel level of a whisper at an intensity of 5.4 × 10-10 watts per

Answer by fractalier(6550)

Okay from
D = 10log (I/10^-12) we plug in the intensity I we are given to find D...
D = 10 * log(5.4 × 10^-10 / 10^-12)
D = 10 * log(5.4 x 10^2)
D = 10 * 2.732
D = 27.32 decibels